Positronium consists of electron and positron orbiting round their common centre of mass. Calculate the value of Rydberg constant for the system?
 

Grade:12th pass

1 Answers

Aziz Alam IIT Roorkee
9 years ago
In Bohr model, for the calculation of Rydberg constant, the mass of an atom is considered to be very large in comparison to electron. The Rydberg constant, hence is denoted as :
R0= 1.097 x 10-7m-1
We know for a fact that that is not the case, .i.e.,nucleus has a finite mass.Hence the Rydberg constant for any atom is given as:
RM=R0(1-me/M) : me=mass of electron, M=Total mass of the atom.
In case of positronium M=2me. Therefore, Rm=R0/2


Thanks &Regards
Aziz Alam
IIT Roorkee
askIITians Faculty